### v1

# MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM

Abstract
Large language models (LLMs) have advanced code generation from single-function tasks to competitive-programming problems, but existing multi-agent solutions either rely on costly large-scale (>30B>30B) models or collapse when downsized to small open-source models. We present MapCoder-Lite, which upgrades a single 7B model into four role-specialised agents—retriever, planner, coder, and debugger—using only rank-32, role-specific LoRA adapters ($3\%$ extra parameters). Three lightweight techniques make this possible: (i) trajectory distillation from strong LLMs fixes format fragility in retrieval and debugging, (ii) supervisor-guided correction strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation. Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, and closes to within six points of a 32B baseline while cutting GPU memory and token-generation time by $4\times$. These results demonstrate that careful agent-wise fine-tuning unleashes high-quality multi-agent coding on a small language model.
MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM
Woongkyu Lee1   Junhee Cho2   Jungwook Choi1thanks: Corresponding author. 1Hanyang University  2Samsung SDS {lwghanyang, choij}@hanyang.ac.kr junhee.cho@samsung.com
1 Introduction
LLMs have revolutionized code synthesis, achieving near-perfect accuracy on function-level tasks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). Research has now shifted to competitive programming, which demands efficient algorithms, robust implementation, and resilience against hidden test cases. Benchmarks like CodeElo Quan et al. (2025) highlight significant challenges that have spurred the development of massive models, such as OpenAI’s o3 OpenAI et al. (2025) and the 671 B-parameter DeepSeek-V3 DeepSeek-AI et al. (2025b), which dominate leaderboards despite their computational costs and proprietary nature.
To meet the cognitive breadth of competitive programming, recent research has shifted from monolithic prompting to multi-agent code-generation frameworks that allocate distinct roles to specialised language-model agents Islam et al. (2024). These systems separate problem solving into discrete stages—algorithm selection, planning, coding, and debugging—enabling focused expertise and structured information exchange between agents Huang et al. (2024); Hong et al. (2024); Islam et al. (2024, 2025). MapCoder Islam et al. (2024) (Fig. 1) exemplifies this approach with a pipeline where specialized agents handle different aspects of problem solving, narrowing the gap between open-source and proprietary solutions.
However, multi-agent systems present two significant challenges. First, they are computationally intensive when using large models. Our MapCoder implementation with Qwen-2.5-32B on the CodeContests benchmark required 3,000 forward passes, each involving 70 GB of parameter loading, resulting in 27.5 hours of runtime on a single A100 GPU—far exceeding the cost of single-prompt approaches. Second, downsizing the agents to 7B models leads to a sharp decline in reliability (see Fig. 2): (i) format adherence breaks down, with frequent XML schema violations that halt pipeline execution; and (ii) role-specific performance suffers, yielding brittle algorithms, incorrect code, and ineffective debugging. Achieving a balance between competitive performance and computational efficiency remains a key open problem.
Figure 1: Overview of the MapCoder system. Given a natural language problem, the retrieval agent fetches relevant algorithmic knowledge, followed by the planning agent generating a solution plan. The coding agent implements the plan, and the debugging agent iteratively refines the code based on test outcomes.
We address these challenges with MapCoder-Lite, the first multi-agent coding framework that drives a single 7B backbone–extended only by lightweight, role-specific adapters–to performance near that of 32B systems. MapCoder-Lite is built on three components:
  • Trajectory distillation from strong LLMs (Sec. 3.1): For retrieval and debugging roles—where small models often fail—we harvest high-quality traces from Qwen-32B and DeepSeek-V3, filtered through execution tests to ensure format correctness and semantic accuracy.
  • Supervisor-aided cross-agent refinement (Sec. 3.2): For planning and coding roles, we employ a supervisor model that identifies failures, provides targeted feedback, and regenerates problematic steps, creating high-quality training examples without human annotation.
  • Memory-efficient LoRA specialization (Sec. 3.3): All agents share a frozen Qwen-2.5-7B backbone with role-specific low-rank adapters (LoRA Hu et al. (2021)), adding <3% of the parameters of a 7B model while enabling independent fine-tuning.
We conducted a comprehensive evaluation on three representative competitive-programming suites—xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, and narrows the gap to a 32B MapCoder to within six points, all while cutting GPU memory and token generation time(time per output token) by $4\times$. These results demonstrate that careful agent-wise fine-tuning can unlock high-quality multi-agent code generation on small language models.
2 Background
2.1 LLMs for Competitive Programming
Large language models have already revolutionised function-level code generation: specialised variants such as Codex, Qwen-2.5-Coder, DeepSeek-Coder, and CodeLlama routinely achieve near-perfect scores on HumanEval, MBPP, and LiveCodeBench Chen et al. (2021); Hui et al. (2024); Guo et al. (2024); Rozière et al. (2024); Jain et al. (2024). But recent community attention has shifted toward the far more demanding arena of competitive programming, where a model must translate a prose problem statement into a complete, time- and memory-efficient program that solves all hidden test cases. Datasets such as APPS Hendrycks et al. (2021), CodeContests Li et al. (2022), and xCodeEval Khan et al. (2024) embody this shift: they assess full-program synthesis, algorithm selection, and constraint reasoning under the stringent standards of contest judging. Progress on these benchmarks now defines the “state of the art” for code-capable LLMs OpenAI et al. (2025); DeepSeek-AI et al. (2025a), but their difficulty also exposes the limits of single-agent prompting, motivating the multi-agent systems discussed next.
2.2 Multi-Agent Code Generation
As competitive programming problems grow exceptionally in linguistic length and algorithmic subtlety, it is not uncommon for a single LLM prompt to miss hidden edge cases or violate the constraints of the competition. Recent work, therefore, turns to multi-agent pipelines, in which a set of specialised LLM agents collaborate under an explicit workflow Huang et al. (2024); Pan et al. (2025). Among these systems, MapCoder Islam et al. (2024) stands out with its attractive design that performs well in APPS and CodeContests. Islam et al. (2024).
Fig. 1 illustrates MapCoder’s four-stage coding pipeline. A retrieval agent first queries a private corpus of textbook algorithms and returns a short, schema-constrained XML snippet that explains a candidate approach. Next, a planning agent expands this snippet into several step-wise execution plans, and a confidence score is generated for each plan using XML format. Plans are sorted by confidence, and the top one is passed to a coding agent that writes full source code in the required language. Finally, a debugging agent compiles, executes, and patches the program until it clears all unit tests; if it fails, the loop backtracks to the next plan.
Role specialisation makes the pipeline modular: each agent faces a simpler, well-scoped sub-task, so errors can be isolated and corrected step-by-step, yielding higher overall accuracy than a monolithic prompt can achieve. The price of this modularity, however, is that the underlying LLM must possess enough general capability to perform four very different roles—retrieving algorithms, reasoning about plans, writing code, and fixing bugs—without losing coherence across stages.
2.3 Task-Specific Fine-Tuning
MapCoder relies on large-scale (>30B>30B), general-purpose LLMs at every stage, trading speed and memory for accuracy. Because each agent tackles a well-defined sub-task—retrieving algorithms, planning steps, writing code, or fixing bugs—a natural question is whether a single 7 B backbone can match MapCoder if it is fine-tuned per role. Fine-tuning smaller models has proved effective in other multi-agent settings Zhao et al. (2025); Liang et al. (2025); Shen et al. (2024), and several studies have enhanced code-related abilities through supervised updates or parameter-efficient adapters Fan et al. (2025); Tsai et al. (2024); Yu et al. (2024); Jiang et al. (2024a). To our knowledge, however, no work has comprehensively studied how far role-aligned fine-tuning can push a small LLM inside a multi-agent pipeline for competitive programming.
Model Metric Cost
Number of problems 165
Total runtime 27.53 hours
Qwen32B Input tokens 5.08M tokens
Output tokens 1.60M tokens
Number of API calls 3,095 calls
Table 1: Resource cost of MapCoder using Qwen2.5-32B-Instruct on the CodeContests benchmark.
Figure 2: Representative failure cases in 7B-scale models. (a) Format following error in the retrieval agent: the model fails to generate responses in the required XML structure, leading to format mismatches and processing errors. (b) Incomplete reasoning in the planning agent: a key step is omitted from the solution plan, resulting in failure to fully cover possible cases.
2.4 Challenges
We identify key challenges in multi-agent code generation based on our implementation of MapCoder.
High Cost with Large Models. We evaluated MapCoder using the Qwen2.5-32B-Instruct model on the CodeContests benchmark, which comprises 165 competitive programming problems. As shown in Table 1, the system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls. This substantial resource usage stems from MapCoder’s multi-agent design involving four distinct agents and multiple iterations for planning and debugging. These results highlight the heavy runtime and memory burden of using large-scale language models (LMs) throughout the pipeline. We hypothesize that replacing each agent with a fine-tuned small LM can significantly reduce token-generation time and GPU memory usage, even with similar token and API call counts.
Failure Cases with Small Models. Fig. 2 illustrates typical failures when downsizing MapCoder’s model to a 7B scale.
  • Format Following Failures: Multi-agent systems often rely on structured outputs in predefined formats. For instance, in MapCoder, the retrieval and planning agents must return XML-formatted responses with tags like <root>, <algorithm> and <confidence>. In Fig. 2(a), the 7B model inserts an extraneous <description> tag and omits the closing </root> tag, violating the expected format. Such structural errors disrupt downstream processing, as subsequent agents or parsers depend on well-formed inputs. These issues are particularly critical in multi-agent pipelines, where format consistency ensures reliable coordination.
  • Low Role Performance: Small models often struggle with role-specific tasks due to limited capacity. As shown in Fig. 2(b), the 32B model’s planning agent enumerates all relevant cases, demonstrating robust and comprehensive reasoning. In contrast, the 7B model omits a key sub-case involving the combination of multiple ticket types. Incomplete planning like this can yield incorrect solutions, even if individual steps seem plausible. Similar shortcomings appear in other roles: the coding agent may generate incorrect logic or indexing, while the debugging agent frequently fails to identify and correct bugs, leading to repeated failures across iterations.
To mitigate these issues, we propose task-specific fine-tuning to improve both format compliance and role execution, thereby enhancing overall code generation accuracy.
3 Methodology
To replace MapCoder’s 32B backbone with a memory-efficient 7B model while maintaining accuracy, we introduce a role-aligned supervised fine-tuning (SFT) pipeline that tailors one shared 7B backbone to four agents—retrieval, planning, coding, and debugging—via lightweight LoRA adapters. These changes present three specific challenges:
  • Role-specific data scarcity: Public code-generation datasets provide ground-truth programs but lack datasets for retrieval queries, step-wise plans, and debug traces, hindering agent-wise supervision.
  • Cross-agent consistency: Naïvely fine-tuning each agent in isolation overlooks inter-agent dependencies; a fix applied to one step can still fail downstream.
  • Parameter budget inflation: Even if fine-tuning recovers agent-specific capability, storing four independent 7B checkpoints nullifies memory savings, approaching the original 32B footprint.
We address (i) by distilling high-quality trajectories from a strong LLM for retrieval & debugging, tackle (ii) with a supervisor model that diagnoses end-to-end failures and regenerates only the faulty segment for planning & coding, and solve (iii) by sharing a single 7B backbone augmented with rank-32 LoRAs—adding $3\%$ parameters while preserving MapCoder’s modularity.
Figure 3: Illustration of trajectory construction for retrieval and debugging datasets. (a) When a strong LLM generates a complete solution that passes unit tests, the retrieval agent’s input-output pair is extracted as a training example. (b) To collect debugging data, a 7B model is used for planning and coding, and the strong LLM is used for debugging when the initial code fails. If the revised output passes, the debugging trajectory is added to the dataset.
3.1 Strong LLM for Retrieval and Debugging
Observations. Two weaknesses repeatedly surface when the 32B backbone is replaced with a 7B model. First, the retrieval agent often violates the mandated XML schema or returns partial algorithm descriptions, so downstream agents cannot even parse its output. Second, the debugging agent fails to patch seemingly simple boundary-case bugs, leaving many programs one step short of success. Experiments with stronger LLMs such as Qwen2.5-32B-Instruct and DeepSeek-V3 show that these larger LLMs seldom make such mistakes, indicating that their trajectories can serve as high-quality supervision for smaller LLMs.
Challenges. Effective role-specific fine-tuning is hampered by two data gaps. (i) Public code-generation benchmarks provide only problem–solution pairs and contain no retrieval queries, step-wise plans, or debug traces; thus, no off-the-shelf corpus targets the intermediate process of the multi-agent coding pipeline. (ii) Even when synthetic trajectories are generated, many are noisy and fail the unit tests, so naïvely training on them would propagate errors rather than correct them.
Methods. Fig. 3 shows the proposed trajectory construction method. Our data pipeline begins by asking strong LLMs—Qwen2.5-32B-Instruct or DeepSeek-V3—to solve each coding task while explicitly printing the intermediate artefacts (i.e. trajectories) obtained from each MapCoder role. In one shot, the model emits an algorithm description for the retrieval agent, a step-wise plan for the planning agent, source code for the coding agent, and a patched version for the debugging agent. Because the same LLM generates the entire trajectory, the four outputs are mutually consistent and naturally aligned with their respective prompts.
To convert these raw trajectories into reliable supervision, we exploit the fact that MapCoder’s final objective—passing all the unit tests—is objectively verifiable. After executing the LLM-produced code, we keep the complete trajectory only if the program passes. This pass‐based filtering eliminates noisy or partially correct traces without any manual inspection: a failure at test time almost always implies that at least one upstream artefact (faulty retrieval cue, flawed plan, or ineffective patch) is wrong, so discarding the whole sequence prevents error propagation during fine-tuning. The surviving set—about 2,200 retrieval examples and 4,200 debugging examples from APPS, xCodeEval, and CodeContests—forms a high-quality, role-aligned corpus that directly trains rank-32 LoRA on the shared 7B backbone. Empirically, models tuned on this filtered dataset regain XML-schema fidelity and bug-repair accuracy comparable to the original 32 B system while retaining the memory footprint advantages of a 7B model (see Table 2 for resulting accuracy enhancement).
Figure 4: Supervisor-aided data collection pipeline. When the final output fails, the supervisor inspects the full trajectory (including algorithm, plan, code, and test result), identifies the responsible agent, and provides targeted feedback to revise its output. If the revised result passes, the updated trajectory is added to the fine-tuning dataset.
Retrieval Planning / Coding xCodeEval
Dataset
Pass
Filtering
Dataset
Pass
Filtering
Accuracy (%)
11.32
Strong 16.04
Strong 16.98
Strong Strong 18.87
Strong Supervisor 22.64
Table 2: Ablation study on data source and pass-based filtering for training retrieval, planning, and coding agents on xCodeEval. “Strong” refers to data generated by a large LM (e.g., Qwen2.5-32B, DeepSeek-V3), and “Supervisor” denotes data refined by a supervisor model after failure analysis.
3.2 Supervisor-Guided Planning and Coding
Observations. Even after the role-specific fine-tuning on strong LLM trajectories, the 7B planning agent still emits fragmentary execution flows, omitting edge cases, misordering steps, or drifting from the prescribed outline. Also, the coding agent continues to generate programs that compile yet collapse under the full unit-test suite. These local errors propagate: an incomplete plan starves the coder of crucial hints, and a superficially correct program deceives the debugger until final evaluation. Table 2 confirms that this trajectory-only adaptation closes part of the gap but suffers from a noticeable accuracy loss. In contrast, Qwen2.5-32B-Instruct and DeepSeek-V3 produce structurally sound plans and largely correct first-pass implementations, indicating that stronger models possess global cues that the smaller model has not yet internalised.
Challenges. Despite the availability of agent-wise strong-model trajectories, two obstacles remain for high-quality fine-tuning. First, a locally tuned agent lacks global awareness: improving its own output does not guarantee end-to-end success unless that output is judged in context. Second, the capacity gap hinders direct imitation: a 7B model exposed to large-scale models’ plans and coding results may underfit or copy surface patterns without learning the deeper reasoning (phenomenon called capacity gap Bansal et al. (2024); Xu et al. (2025)).
Methods. We propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model. As shown in Fig. 4, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory. If the resulting program fails its unit tests, the entire trajectory is forwarded to a high-capacity supervisor LLM (DeepSeek-V3). The supervisor analyses the trace, pinpoints the agent primarily responsible for the failure, and issues concise, role-specific feedback, and only that agent regenerates its output. The revised trajectory is re-tested; once all tests pass, the final plan–code pair is added to the fine-tuning corpus.
Supervisor-guided refinement is applied selectively to trajectories where the strong LLM succeeds but the 7B model fails, minimizing generation overhead. The supervisor operates only during data generation, keeping inference lightweight. We store just the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime. Crucially, every example in this corpus is produced and execution-validated by the 7B model, eliminating concerns about a capacity gap. Iterating over thousands of problems yields a dataset that aligns planners and coders with end-to-end success. Table 2 confirms that this strategy delivers a substantial accuracy boost.
FT Method
Trainable
Parameters
Required Memory
for Training (GB)
Accuracy
FFT 7615.62M 45.69 18.87
LoRA 20.19M 15.35 22.64
Table 3: The number of trainable parameters, required memory for training, and xCodeEval accuracy of full fine-tuning (FFT) and LoRA methods with MapCoder-Lite 7B. Accuracy is measured only up to the coding stage, without debugging.
3.3 Multi-Agent LoRA Fine-Tuning
To enable role-specific specialization in our multi-agent framework, we adopt an agent-wise LoRA approach. All four agents—retrieval, planning, coding, and debugging—share a frozen 7B base model, each paired with its own lightweight LoRA. These adapters are fine-tuned independently using pass-based-filtered and supervisor-guided trajectory distillation. This setup significantly reduces GPU memory usage during fine-tuning and supports modular training and deployment, allowing each agent to be updated without duplicating the full model. As shown in Table 3, LoRA significantly reduces the number of trainable parameters and memory footprint, while achieving even higher accuracy than full fine-tuning. We attribute this accuracy gain in part to LoRA’s implicit regularization, which steers the model toward role-specific behaviors without overwriting core pretrained knowledge—a finding consistent with prior work Hu et al. (2021); Houlsby et al. (2019).
4 Experimental Settings
Method xCodeEval (106) APPS (150) CodeContests (165)
[…]
Direct 18.87 20 4.00 6 1.21 2
CoT 5.66 6 4.67 7 3.64 6
Self-planning 9.43 10 1.33 2 3.03 5
Analogical 12.26 13 3.33 5 1.82 3
MapCoder 13.21 14 29 6.00 9 30 6.06 10 59
MapCoder-Lite 28.30 29 0 8.00 12 0 13.33 22 0
Table 4: Performance comparison of different methods using Qwen7B on xCodeEval, APPS, and CodeContests with Pass@1 accuracy (%). The numbers in parentheses indicate the number of problems in each benchmark.
Models. All agent-wise fine-tuning experiments use Qwen2.5-7B-Instruct as the base model. For comparison, we include larger models—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, serving as upper bounds to assess the effectiveness of our lightweight strategy.
Benchmarks. We evaluate models on three competitive programming benchmarks: xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022). For xCodeEval, we use the 106 problems from the compact program synthesis split. For CodeContests, we evaluate all 165 official test problems. We randomly select 150 test-set problems for APPS, following MapCoder Islam et al. (2024). Each benchmark includes natural language problem descriptions with sample I/O and is evaluated using hidden unit tests. We also optionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
Training Configuration. Each agent is fine-tuned using LoRA adapters with rank 32 applied to the linear modules for query, key, value, and output projection. Training is implemented with Hugging Face Transformers and PEFT libraries HuggingFace , using a learning rate of 2e-5, gradient accumulation of 16, and 3 training epochs. Hyperparameters were tuned individually per agent to ensure stable and consistent performance. All training experiments were conducted on a single NVIDIA A100 80GB GPU.
Evaluation Setting. We report pass@1 accuracy as the primary metric, reflecting whether the model’s top-1 output passes all test cases. All outputs are generated using greedy decoding for reproducibility and consistency with baseline evaluation.
5 Experimental Results
5.1 Effectiveness of Agent-wise Fine-Tuning
Table 4 compares the performance of a small LM (Qwen2.5-7B) across various single-agent prompt optimization methods—direct prompting, chain-of-thought (CoT) Wei et al. (2023), self-planning Jiang et al. (2024b), and analogical prompting Yasunaga et al. (2024)—as well as multi-agent approaches: MapCoder and our proposed MapCoder-Lite. Without fine-tuning, MapCoder shows only marginal gains over single-agent methods and even underperforms Direct prompting on xCodeEval (13.21% vs. 18.87%), despite using four specialized agents. This highlights the limitations of small LMs in multi-agent setups, where poor coordination and frequent format violations (as indicated by numerous format failures) prevent reliable end-to-end performance.
In contrast, MapCoder-Lite, with agent-wise fine-tuning, achieves substantial improvements across all benchmarks. On xCodeEval, accuracy more than doubles to 28.30%, outperforming all prompting baselines. Similar gains are seen on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%). Notably, fine-tuning eliminates all format failures, indicating that even small LMs can produce structurally consistent outputs when role-specialized. These results emphasize the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks.
5.2 Comparison with Larger Models
Table 5 compares MapCoder’s performance across backbone sizes from 7B to 32B (Qwen[7–32]B), alongside MapCoder-Lite—our fine-tuned 7B variant (Qwen7B-FT). Qwen7B-FT delivers a substantial accuracy boost over the untuned Qwen7B, matching the performance of Qwen14B and coming within six points of Qwen32B. This is particularly notable given that Qwen7B-FT requires only one-quarter of the GPU memory of Qwen32B, making it viable for deployment on memory-constrained edge devices Karumbunathan (2022). Moreover, the reduced model size yields a proportional speedup in token generation—LLM decoding being memory-bound—achieving a 4× reduction in time-per-output-token (TPOT), as measured using vLLM Kwon et al. (2023). These results highlight that targeted fine-tuning enables small models to achieve competitive performance with dramatically lower cost.
Model xCodeEval APPS CodeContests Memory TPOT
Qwen7B 13.21 6.00 6.06 15.26 12.29
Qwen7B-FT 28.30 8.00 13.33 15.64 12.29
Qwen14B 28.30 10.00 15.76 29.58 23.36
Qwen32B 33.02 13.33 18.18 65.56 45.06
Table 5: Pass@1 accuracy (%) of MapCoder using Qwen models of different scales on xCodeEval, APPS, and CodeContests. Memory usage (GB) and Time Per Output Tokens (ms).
Agent xCodeEval Pass Pass Format
[…]
- - - - 13.21 11 3 29
- - - 18.87 17 3 12
- - 26.42 18 10 0
- 24.53 24 2 0
- 27.36 17 12 0
28.30 24 6 0
Table 6: Ablation study of per-agent fine-tuning contributions in MapCoder on xCodeEval. R, P, C, and D denote Retrieval, Planning, Coding, and Debugging agents, respectively.
Method HumanEval (164 problems) MBPP (397 problems)
[…]
MapCoder 70.73 105 11 14 67.51 244 24 68
MapCoder-Lite 82.93 120 16 1 84.63 305 31 0
Table 7: Evaluation of MapCoder and MapCoder-Lite on HumanEval and MBPP benchmarks.
5.3 Ablation Study
Contribution of Individual Agent Fine-Tuning. Table 6 shows an ablation study evaluating the impact of fine-tuning each agent in MapCoder on xCodeEval. Fine-tuning the retrieval agent alone improves accuracy from 13.21% to 18.87% and reduces format failures from 29 to 12, indicating better structural consistency. Adding the planning agent further boosts accuracy to 26.42% and eliminates format errors entirely, increasing both non-debug passes (17 → 18) and debug-assisted completions (3 → 10). Fine-tuning the coding agent (without debugging) raises non-debug passes to 24 but reduces debug-assisted passes to 2, since MapCoder only invokes the debugger when initial code fails sample I/O tests. Stronger code may bypass debugging entirely, leaving some errors uncaught. Conversely, fine-tuning the debugging agent (without coding) increases accuracy to 27.36%, with 12 debug-assisted completions—showing that an effective debugger can recover from weaker initial code. When all four agents are fine-tuned, overall accuracy peaks at 28.30%, with a strong balance between direct (24) and debug-assisted (6) completions. These results demonstrate that each agent addresses distinct failure modes and that coordinated fine-tuning is essential for maximizing end-to-end performance.
Evaluation on Simpler Coding Tasks. To evaluate generalization beyond competition-level programming, we compare MapCoder and MapCoder-Lite on two function-level benchmarks: HumanEval and MBPP. As shown in Table 7, MapCoder-Lite outperforms the baseline by significantly reducing format-related failures and achieving higher end-to-end accuracy. It improves both coding (increased passes without debugging) and debugging (even greater gains in passes with debugging), indicating enhanced reliability. These results suggest that MapCoder-Lite not only preserves generalization capability but also improves robustness on structurally simpler tasks.
Multi-Agent Tuning with Coder Model. We evaluate whether multi-agent code generation can be enhanced using a code-specialized language model (a.k.a. Coder model), such as Qwen2.5-Coder-7B-Instruct Hui et al. (2024). As shown in Table 8, applying the multi-agent framework (MapCoder) and our fine-tuning strategy (MapCoder-Lite) to the Coder model does improve its performance, demonstrating that even domain-specialized models benefit from structured, role-specific tuning. However, despite these gains, the Coder-based pipeline still lags significantly behind our baseline model, Qwen2.5-7B-Instruct, both before and after fine-tuning. We attribute this gap to the Coder model’s limited general reasoning ability, which is critical for handling the diverse roles in multi-agent workflows, including retrieval, planning, and debugging, in addition to coding. In contrast, the general-purpose Qwen2.5-7B-Instruct model more effectively supports multi-agent coordination, and with our memory-efficient fine-tuning (MapCoder-Lite), achieves even higher end-to-end accuracy.
Model Method Accuracy (%)
Qwen2.5-Coder-7B-Instruct MapCoder 5.66
MapCoder-Lite 15.09
Qwen2.5-7B-Instruct MapCoder 11.32
MapCoder-Lite 22.64
Table 8: Comparison of MapCoder and MapCoder Lite using Qwen2.5-Coder-7B-Instruct. Accuracy is measured only up to the coding stage, without debugging.
5.4 Case Study
To qualitatively assess the impact of agent-wise fine-tuning, we examine representative before-and-after examples for each agent. As illustrated in Appendix B, fine-tuning leads to notable improvements in algorithm retrieval, plan coverage, code correctness, and bug resolution. These examples highlight how each agent benefits from role-specific supervision, supporting the effectiveness of our fine-tuning strategy.
6 Conclusion
We introduced MapCoder-Lite, a multi-agent coding framework built on a 7B model with lightweight LoRA adapters. Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance with lower cost. Our results show that small models, when carefully fine-tuned, can match the reliability and accuracy of larger systems.
Limitations
Our framework relies on distilled trajectories from strong models like DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance. Rather than indicating a ceiling, this gap highlights opportunities to further enhance small models—potentially through architectural extensions, better training objectives, or reinforcement-based learning from distillation feedback.
In addition, the multi-agent structure introduces a wide design space for fine-tuning, as each agent has a distinct role and optimization goal. While we adopted a uniform tuning recipe across agents, this modularity also enables targeted specialization and co-training strategies in future work, suggesting that more adaptive tuning could unlock even greater performance.
Appendix A Details of Supervisor
Fig. 5 and Fig. 6 illustrate the operation of the supervisor used in our data curation pipeline. The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, plan, code), and the result of test case execution.
As shown in Fig. 5, the input prompt explicitly instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent. The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that can be used to re-generate only the faulty component.
Fig. 6 shows a typical response from the supervisor. It correctly attributes the failure to the planning agent, explains why the plan is insufficient, and suggests how it should be modified. This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without requiring human annotation.
Appendix B Improvements After Fine-Tuning
Fig. 789, and 10 illustrate representative improvements observed in the 7B model for the retrieval, planning, coding, and debugging agents, respectively.
In the retrieval example (Fig. 7), the pre-fine-tuned model incorrectly identifies the core algorithm as “Sliding Window” and includes unsupported tags like <description> within the <algorithm> block, failing to conform to the expected XML schema. After fine-tuning, the agent accurately identifies the algorithm (“Counting and Matching Pairs”) and outputs a well-structured XML response with all required tags properly closed.
In the planning example (Fig. 8), the original plan fails to capture key logical conditions (e.g., that the input must be both even and greater than 2). After fine-tuning, the agent successfully generates a plan that handles these conditions explicitly and correctly guides the coding agent.
In the coding example (Fig. 9), the unfine-tuned model incorrectly processes multiple input values on a single line, violating the problem’s input specification. Fine-tuning enables the agent to read inputs line by line, aligning its behavior with the expected input format and improving functional correctness.
Finally, the debugging example (Fig. 10) demonstrates how the unfine-tuned model fails to fix a bug caused by improper input parsing. The fine-tuned debugging agent correctly diagnoses the issue and proposes a revised plan and code that successfully passes all test cases.
These examples highlight the role-specific benefits of fine-tuning, improving not only task-level correctness but also structural and logical alignment across the entire pipeline.
Figure 5: Prompt for the Supervisor.
Figure 6: Example response from the supervisor: identifies the flawed part of the plan and generates feedback.
Figure 7: Improvement in algorithm tutorial and XML formatting by the retrieval agent after fine-tuning. The pre-trained model fails to identify the correct algorithm and generates an ill-formed XML response with an unclosed <root> tag. After fine-tuning, the retrieval agent provides a valid explanation based on matching command pairs and generates a well-structured XML block.
Figure 8: Improvement in conditional logic by the planning agent after fine-tuning. The initial code fails to account for the edge case (w > 2) and only checks for evenness. After fine-tuning, the planning agent correctly generates the full condition required by the problem.
Figure 9: Example of a coding error and its resolution after fine-tuning. The unfine-tuned agent incorrectly reads all input values from a single line, which violates the problem specification. After fine-tuning, the coding agent correctly processes line-separated inputs and computes the result accordingly.
Figure 10: Example of a debugging failure and recovery in the MapCoder pipeline. The initial plan and code fail to address the problem due to incorrect input parsing. After the debugging agent identifies the issue, the revised plan introduces line-by-line input reading, which successfully resolves the error.

### v2

@@
-MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM
+MapCoder-Lite: Distilling Multi-Agent Coding into a Single Small LLM
 Abstract
 Large language models (LLMs) have advanced code generation from single-function tasks to competitive-programming problems, but existing multi-agent solutions either rely on costly large-scale (>30B>30B) models or collapse when downsized to small open-source models.
-We present MapCoder-Lite, which upgrades a single 7B model into four role-specialised agents—retriever, planner, coder, and debugger—using only rank-32, role-specific LoRA adapters ($3\%$ extra parameters).
-Three lightweight techniques make this possible: (i) trajectory distillation from strong LLMs fixes format fragility in retrieval and debugging, (ii) supervisor-guided correction strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation.
-Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, and closes to within six points of a 32B baseline while cutting GPU memory and token-generation time by $4\times$.
+We present MapCoder-Lite, a framework for distilling the complex reasoning of large, multi-agent coding systems into a single 7B model.
+Our contribution is a novel, three-pillar methodology that synergistically generates, refines, and encodes multi-agent knowledge: (i) pass-based trajectory distillation from strong LLMs fixes format fragility in retrieval and reduces failures in debugging, (ii) supervisor-guided correction with global feedback strengthens planning and coding agents, and (iii) agent-wise LoRA fine-tuning delivers memory-efficient specialisation.
+Comprehensive evaluation on xCodeEval, APPS, and CodeContests shows that MapCoder-Lite more than doubles xCodeEval accuracy (13.2% → 28.3%), eliminates all format failures, while reducing GPU memory and token-generation time by $4\times$ compared to a 32B model.
+It also achieves over 10% gains on simpler coding benchmarks, demonstrating broad improvements beyond competitive programming.
 These results demonstrate that careful agent-wise fine-tuning unleashes high-quality multi-agent coding on a small language model.
-MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM
+Our code is publicly available at https://github.com/aiha-lab/MapCoder-Lite.
+MapCoder-Lite: Distilling Multi-Agent Coding into a Single Small LLM
 Woongkyu Lee1   Junhee Cho2   Jungwook Choi1thanks: Corresponding author. 1Hanyang University  2Samsung SDS {lwghanyang, choij}@hanyang.ac.kr junhee.cho@samsung.com
@@
 (2025) and the 671B-parameter DeepSeek-V3 DeepSeek-AI et al.
-(2025b), which dominate leaderboards despite their computational costs and proprietary nature.
-To meet the cognitive breadth of competitive programming, recent research has shifted from monolithic prompting to multi-agent code-generation frameworks that allocate distinct roles to specialised language-model agents Islam et al.
+(2025b).
+These models report strong performance on Codeforces and related benchmarks despite their computational costs and proprietary nature.
+Competitive programming involves algorithmically complex problem solving under strict constraints, making it a difficult setting for language models.
+Single-agent prompting, where one model handles the entire problem-solving process, often falls short Wei et al.
+(2023); Jiang et al.
+(2024b); Yasunaga et al.
 (2024).
-These systems separate problem solving into discrete stages—algorithm selection, planning, coding, and debugging—enabling focused expertise and structured information exchange between agents Huang et al.
+To overcome this, recent work has explored multi-agent code-generation frameworks that split the task into stages and assign each to a dedicated agent, improving end-to-end performance through role specialization Huang et al.
 (2024); Hong et al.
@@
 (2024) (Fig.
-1) exemplifies this approach with a pipeline where specialized agents handle different aspects of problem solving, narrowing the gap between open-source and proprietary solutions.
-However, multi-agent systems present two significant challenges.
-First, they are computationally intensive when using large models.
-Our MapCoder implementation with Qwen-2.5-32B on the CodeContests benchmark required 3,000 forward passes, each involving 70 GB of parameter loading, resulting in 27.5 hours of runtime on a single A100 GPU—far exceeding the cost of single-prompt approaches.
-Second, downsizing the agents to 7B models leads to a sharp decline in reliability (see Fig.
-2): (i) format adherence breaks down, with frequent XML schema violations that halt pipeline execution; and (ii) role-specific performance suffers, yielding brittle algorithms, incorrect code, and ineffective debugging.
-Achieving a balance between competitive performance and computational efficiency remains a key open problem.
+1) exemplifies this approach, coordinating specialized agents throughout the pipeline.
+The effectiveness of multi-agent frameworks typically relies on large-scale (>30B>30B) open models OpenAI et al.
+(2024); Comanici et al.
+(2025), which possess the capacity to perform a wide range of specialized roles required across the multi-agent pipeline.
+Due to their multi-step nature, these frameworks incur significantly higher token usage and API calls than single-agent setups, resulting in increased latency and computational cost.
 Figure 1: Overview of the MapCoder system.
@@
 The coding agent implements the plan, and the debugging agent iteratively refines the code based on test outcomes.
+Their impracticality in resource-constrained settings naturally motivates small-model multi-agent solutions, which align well with the growing industry trend toward on-device AI Gunter et al.
+(2024); Zhang et al.
+(2023).
+In practice, however, deploying small language models (SLMs) under 10B parameters—even those with strong coding abilities Qwen et al.
+(2025)—often yields limited accuracy gains compared to single-model direct prompting.
+This performance drop stems from SLMs’ difficulty in following the structured formats (e.g., XML) required for multi-agent communication Xia et al.
+(2024) and their limited capacity to support the complex reasoning needed across agent roles, leading to failures in retrieval, planning, or debugging.
+To make SLM-based multi-agent systems effective, fine-tuning becomes a necessary step.
+However, this approach poses three key obstacles.
+First, existing code datasets do not provide intermediate artefacts aligned with the roles in a multi-agent pipeline.
+Second, fine-tuning each agent independently fails to account for inter-agent dependencies.
+As illustrated in Fig.
+1, later stages rely on the outputs of earlier agents, so errors in one stage can propagate and mislead downstream components.
+Third, training separate models for each agent increases GPU memory consumption at inference time, undermining the efficiency benefits of using small models in the first place.
 We address these challenges with MapCoder-Lite, the first multi-agent coding framework that drives a single 7B backbone–extended only by lightweight, role-specific adapters–to performance near that of 32B systems.
 MapCoder-Lite is built on three components:
-  • Trajectory distillation from strong LLMs (Sec.
-3.1): For retrieval and debugging roles—where small models often fail—we harvest high-quality traces from Qwen-32B and DeepSeek-V3, filtered through execution tests to ensure format correctness and semantic accuracy.
+  • Pass-based trajectory distillation from strong LLMs (Sect.
+5.1): To address the lack of role-specific training data, we collect trajectories from strong LLMs.
+However, fine-tuning on outputs that are only locally valid often fails to yield correct final solutions.
+We overcome this by implementing pass-based filtering that exclusively retains trajectories whose final code passes all unit tests.
   • Supervisor-aided cross-agent refinement (Sec.
-3.2): For planning and coding roles, we employ a supervisor model that identifies failures, provides targeted feedback, and regenerates problematic steps, creating high-quality training examples without human annotation.
+5.2): We employ a supervisor model that analyzes the full trajectory generated by the small model to detect cross-agent failure patterns and regenerate the responsible agent’s output, guiding the model toward global success and helping bridge the capacity gap.
   • Memory-efficient LoRA specialization (Sec.
-3.3): All agents share a frozen Qwen-2.5-7B backbone with role-specific low-rank adapters (LoRA Hu et al.
-(2021)), adding <3% of the parameters of a 7B model while enabling independent fine-tuning.
+5.3): We show that in complex multi-agent settings, LoRA Hu et al.
+(2021) achieves better accuracy and efficiency than full fine-tuning.
+This enables all agents to share a frozen Qwen2.5-7B backbone with lightweight, role-specific adapters, adding under 3% extra parameters.
 We conducted a comprehensive evaluation on three representative competitive-programming suites—xCodeEval Khan et al.
@@
 (2021), and CodeContests Li et al.
-(2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, and narrows the gap to a 32B MapCoder to within six points, all while cutting GPU memory and token generation time(time per output token) by $4\times$.
+(2022)—and found that MapCoder-Lite leverages trajectory distillation, supervisor-guided cross-agent refinement, and rank-32 LoRA specialisation to boost xCodeEval accuracy from 13.2% to 28.3%, eliminates every XML-schema failure, all while cutting GPU memory and token generation time(time per output token) by $4\times$.
 These results demonstrate that careful agent-wise fine-tuning can unlock high-quality multi-agent code generation on small language models.
-2 Background
+2 Related Work
+Figure 2: Representative failure cases of the 7B-scale model across all agents.
+(a) Retrieval: invalid XML format and incorrect algorithm.
+(b) Planning: missing key step.
+(c) Coding: misinterpreted input specification.
+(d) Debugging: persistent unresolved error.
+Detailed illustrations are provided in Appendix E.
 2.1 LLMs for Competitive Programming
-Large language models have already revolutionised function-level code generation: specialised variants such as Codex, Qwen-2.5-Coder, DeepSeek-Coder, and CodeLlama routinely achieve near-perfect scores on HumanEval, MBPP, and LiveCodeBench Chen et al.
-(2021); Hui et al.
+LLMs have shown strong performance on function-level code generation tasks Hui et al.
 (2024); Guo et al.
-(2024); Rozière et al.
-(2024); Jain et al.
+(2024); rozière2024codellama, achieving near-perfect scores on HumanEval Chen et al.
+(2021) and MBPP Austin et al.
 (2021).
-But recent community attention has shifted toward the far more demanding arena of competitive programming, where a model must translate a prose problem statement into a complete, time- and memory-efficient program that solves all hidden test cases.
-Datasets such as APPS Hendrycks et al.
+Recent work shifts to the harder setting of competitive programming, which requires generating full, efficient programs that pass hidden tests.
+Benchmarks like APPS Hendrycks et al.
 (2021), CodeContests Li et al.
 (2022), and xCodeEval Khan et al.
-(2024) embody this shift: they assess full-program synthesis, algorithm selection, and constraint reasoning under the stringent standards of contest judging.
-Progress on these benchmarks now defines the “state of the art” for code-capable LLMs OpenAI et al.
+(2024) capture this challenge and now define the state of the art OpenAI et al.
 (2025); DeepSeek-AI et al.
-(2025a), but their difficulty also exposes the limits of single-agent prompting, motivating the multi-agent systems discussed next.
+(2025a), motivating the multi-agent systems that follow.
 2.2 Multi-Agent Code Generation
-As competitive programming problems grow exceptionally in linguistic length and algorithmic subtlety, it is not uncommon for a single LLM prompt to miss hidden edge cases or violate the constraints of the competition.
-Recent work, therefore, turns to multi-agent pipelines, in which a set of specialised LLM agents collaborate under an explicit workflow Huang et al.
+Recent work has proposed multi-agent pipelines to better handle the linguistic complexity and algorithmic subtlety of competitive programming, which often cause single prompts to miss edge cases or violate constraints Huang et al.
 (2024); Pan et al.
 (2025).
-Among these systems, MapCoder Islam et al.
-(2024) stands out with its attractive design that performs well in APPS and CodeContests.
-Islam et al.
-(2024).
-Fig.
-1 illustrates MapCoder’s four-stage coding pipeline.
-A retrieval agent first queries a private corpus of textbook algorithms and returns a short, schema-constrained XML snippet that explains a candidate approach.
-Next, a planning agent expands this snippet into several step-wise execution plans, and a confidence score is generated for each plan using XML format.
-Plans are sorted by confidence, and the top one is passed to a coding agent that writes full source code in the required language.
-Finally, a debugging agent compiles, executes, and patches the program until it clears all unit tests; if it fails, the loop backtracks to the next plan.
-Role specialisation makes the pipeline modular: each agent faces a simpler, well-scoped sub-task, so errors can be isolated and corrected step-by-step, yielding higher overall accuracy than a monolithic prompt can achieve.
-The price of this modularity, however, is that the underlying LLM must possess enough general capability to perform four very different roles—retrieving algorithms, reasoning about plans, writing code, and fixing bugs—without losing coherence across stages.
+Among these, MapCoderIslam et al.
+(2024) adopts a four-stage pipeline (Fig.1) that performs well on APPS and CodeContests.
+A retrieval agent first selects an algorithm from a private corpus and returns a schema-constrained XML snippet.
+A planning agent then expands this into step-wise plans with confidence scores.
+The top plan is passed to a coding agent, which generates the full source code.
+Finally, a debugging agent tests and patches the code until it passes all unit tests, backtracking to alternative plans if needed.
+This modular design improves accuracy via role specialization, but also demands broad capability from the underlying LLM across all stages.
 2.3 Task-Specific Fine-Tuning
-MapCoder relies on large-scale (>30B>30B), general-purpose LLMs at every stage, trading speed and memory for accuracy.
-Because each agent tackles a well-defined sub-task—retrieving algorithms, planning steps, writing code, or fixing bugs—a natural question is whether a single 7 B backbone can match MapCoder if it is fine-tuned per role.
-Fine-tuning smaller models has proved effective in other multi-agent settings Zhao et al.
+MapCoder uses strong LLMs at each stage, trading efficiency for accuracy.
+A natural question is whether a single small model (<10B), fine-tuned per role, can achieve similar results.
+While prior work has shown the value of fine-tuning in multi-agent Zhao et al.
 (2025); Liang et al.
 (2025); Shen et al.
-(2024), and several studies have enhanced code-related abilities through supervised updates or parameter-efficient adapters Fan et al.
+(2024) and code tasks Fan et al.
 (2025); Tsai et al.
@@
 (2024); Jiang et al.
-(2024a).
+(2024a), no study has systematically explored role-aligned fine-tuning for small LLMs in competitive programming pipelines.
 To our knowledge, however, no work has comprehensively studied how far role-aligned fine-tuning can push a small LLM inside a multi-agent pipeline for competitive programming.
-Model Metric Cost
-Number of problems 165
-Total runtime 27.53 hours
-Qwen32B Input tokens 5.08M tokens
-Output tokens 1.60M tokens
-Number of API calls 3,095 calls
-Table 1: Resource cost of MapCoder using Qwen2.5-32B-Instruct on the CodeContests benchmark.
-Figure 2: Representative failure cases in 7B-scale models.
-(a) Format following error in the retrieval agent: the model fails to generate responses in the required XML structure, leading to format mismatches and processing errors.
-(b) Incomplete reasoning in the planning agent: a key step is omitted from the solution plan, resulting in failure to fully cover possible cases.
-2.4 Challenges
-We identify key challenges in multi-agent code generation based on our implementation of MapCoder.
-High Cost with Large Models.
+3 Analysis of Multi-Agent Limitations
+3.1 High Cost with Large Models
 We evaluated MapCoder using the Qwen2.5-32B-Instruct model on the CodeContests benchmark, which comprises 165 competitive programming problems.
-As shown in Table 1, the system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls.
+The system required 27.53 hours of runtime, processed approximately 5.08 million input tokens and 1.60 million output tokens, and made 3,095 API calls.
 This substantial resource usage stems from MapCoder’s multi-agent design involving four distinct agents and multiple iterations for planning and debugging.
-These results highlight the heavy runtime and memory burden of using large-scale language models (LMs) throughout the pipeline.
-We hypothesize that replacing each agent with a fine-tuned small LM can significantly reduce token-generation time and GPU memory usage, even with similar token and API call counts.
-Failure Cases with Small Models. Fig. 2 illustrates typical failures when downsizing MapCoder’s model to a 7B scale.
-  • Format Following Failures: Multi-agent systems often rely on structured outputs in predefined formats.
-For instance, in MapCoder, the retrieval and planning agents must return XML-formatted responses with tags like <root>, <algorithm> and <confidence>.
-In Fig.
-2(a), the 7B model inserts an extraneous <description> tag and omits the closing </root> tag, violating the expected format.
-Such structural errors disrupt downstream processing, as subsequent agents or parsers depend on well-formed inputs.
-These issues are particularly critical in multi-agent pipelines, where format consistency ensures reliable coordination.
-  • Low Role Performance: Small models often struggle with role-specific tasks due to limited capacity.
-As shown in Fig.
-2(b), the 32B model’s planning agent enumerates all relevant cases, demonstrating robust and comprehensive reasoning.
-In contrast, the 7B model omits a key sub-case involving the combination of multiple ticket types.
-Incomplete planning like this can yield incorrect solutions, even if individual steps seem plausible.
-Similar shortcomings appear in other roles: the coding agent may generate incorrect logic or indexing, while the debugging agent frequently fails to identify and correct bugs, leading to repeated failures across iterations.
-To mitigate these issues, we propose task-specific fine-tuning to improve both format compliance and role execution, thereby enhancing overall code generation accuracy.
-3 Methodology
-To replace MapCoder’s 32B backbone with a memory-efficient 7B model while maintaining accuracy, we introduce a role-aligned supervised fine-tuning (SFT) pipeline that tailors one shared 7B backbone to four agents—retrieval, planning, coding, and debugging—via lightweight LoRA adapters.
-These changes present three specific challenges:
-  • Role-specific data scarcity: Public code-generation datasets provide ground-truth programs but lack datasets for retrieval queries, step-wise plans, and debug traces, hindering agent-wise supervision.
-  • Cross-agent consistency: Naïvely fine-tuning each agent in isolation overlooks inter-agent dependencies; a fix applied to one step can still fail downstream.
-  • Parameter budget inflation: Even if fine-tuning recovers agent-specific capability, storing four independent 7B checkpoints nullifies memory savings, approaching the original 32B footprint.
-We address (i) by distilling high-quality trajectories from a strong LLM for retrieval & debugging, tackle (ii) with a supervisor model that diagnoses end-to-end failures and regenerates only the faulty segment for planning & coding, and solve (iii) by sharing a single 7B backbone augmented with rank-32 LoRAs—adding $3\%$ parameters while preserving MapCoder’s modularity.
+These results highlight the heavy runtime and memory burden of using large-scale language models throughout the pipeline.
+We therefore hypothesize that replacing each agent with a fine-tuned SLM can substantially reduce token-generation time and GPU memory usage, even under comparable token and API call counts.
+3.2 Failure Cases of Small Models
+Format Following Failures.
+Multi-agent systems often rely on structured outputs in predefined formats Yang et al.
+(2025); Tang et al.
+(2025).
+In MapCoder, for example, retrieval and planning agents are required to produce XML-formatted responses such as <root>, <algorithm>, and <confidence> for downstream parsing.
+However, small models frequently violate these schemas—omitting or misplacing tags—which disrupts subsequent processing and halts pipeline execution (Fig.
+2a).
+This issue is amplified by the weaker format-following ability of open-source SLMs compared to proprietary LLMs Xia et al.
+(2024), underscoring the importance of strict structural adherence in multi-agent workflows.
+Low Role Performance.
+Small models often struggle with role-specific tasks due to limited capacity.
+Fig.
+2 illustrates representative failures across agents: (a) the retrieval agent produces incorrect or misleading algorithm descriptions, corrupting shared context for all downstream stages; (b) the planning agent outputs superficially correct but incomplete plans, often missing subtle edge cases; (c) the coding agent introduces logical or I/O errors even when given valid plans, yielding incorrect or unexecutable code; and (d) the debugging agent fails to detect or fix simple bugs, resulting in repeated ineffective patches.
+Together, these failures indicate that, without stronger supervision or greater capacity, individual agents act as bottlenecks that undermine the reliability of the entire pipeline.
+4 Challenges
+The multi-agent approach provides a parameter-free mechanism for orchestrating specialized roles via prompting, yet SLMs consistently underperform in this zero-shot setting.
+This suggests that prompt-based assignment alone is insufficient for SLMs, necessitating a shift toward parameter-driven optimization Du et al.
+(2025).
+However, effectively fine-tuning SLMs within a multi-agent framework presents several non-trivial challenges.
+4.1 Lack of Role-Specific Training Data.
+Multi-agent fine-tuning requires high-quality intermediate supervision tailored to each agent’s role.
+Existing approaches rely either on reconstructing intermediate signals from public datasets Shen et al.
+(2024) or on self-collection using SLM-generated outputs Liang et al.
+(2025).
+However, public code benchmarks are not designed for multi-agent pipelines and lack role-aligned artefacts, while self-collected data from small models is often malformed or incomplete due to limited capacity and task complexity.
+This underscores the need to generate new, role-specific supervision tailored to multi-agent training.
 Figure 3: Illustration of trajectory construction for retrieval and debugging datasets.
@@
 If the revised output passes, the debugging trajectory is added to the dataset.
+4.2 Limited Global Awareness.
+In multi-agent workflows, information flows sequentially, making end-to-end success reliant on the coherence of intermediate outputs across stages.
+Yet agents trained in isolation lack awareness of these dependencies, leading to local errors that propagate and compromise the final outcome.
+For instance, when the final code fails, it is often unclear whether the root cause lies in a flawed plan from the planning agent or in an incorrect implementation by the coding agent, even if each agent appears to perform its role plausibly.
+Such failures show that success depends not only on individual competence but also on aligned interactions.
+Without global awareness, even well-tuned agents may fall short.
+4.3 Inefficiency of Full Model Fine-Tuning
+Previous multi-agent fine-tuning approaches typically rely on OpenAI’s fine-tuning API Zhao et al.
+(2025) or full-parameter tuning for each agent Shen et al.
+(2024); Zeng et al.
+(2023).
+However, full fine-tuning scales memory usage linearly with the number of agents, negating one of the primary advantages of SLMs—their efficiency in memory and deployment.
+Furthermore, full fine-tuning does not necessarily guarantee superior performance over parameter-efficient methods such as LoRA Hu et al.
+(2021).
+5 Methodology
+To overcome the limitations of applying multi-agent code generation to SLMs, we propose a role-aligned supervised fine-tuning pipeline that equips a single 7B backbone with specialized behavior for four agents via lightweight LoRA adapters.
+This design directly addresses the three main challenges identified in our analysis:
+  • Lack of role-specific data.
+To compensate for the absence of intermediate supervision, we distill high-quality trajectories from strong LLMs, using execution results to retain only clean, agent-specific samples (Sec.
+5.1).
+  • Limited global awareness.
+To maintain cross-agent consistency, we introduce a supervisor-guided refinement mechanism that detects failures and regenerates only the faulty component, yielding coherent, context-aware training data (Sec.
+5.2).
+  • Inefficiency of full-model fine-tuning.
+We address parameter and memory overhead by applying rank-32 LoRA adapters to a shared frozen 7B backbone, enabling agent-wise specialization with minimal additional cost (Sec.
+5.3).
+Together, these techniques allow us to retain the advantages of the multi-agent approach while making it viable for deployment on open-source, resource-efficient language models.
 5.1 Strong LLM for Retrieval and Debugging
-Observations.
-Two weaknesses repeatedly surface when the 32B backbone is replaced with a 7B model.
-First, the retrieval agent often violates the mandated XML schema or returns partial algorithm descriptions, so downstream agents cannot even parse its output.
-Second, the debugging agent fails to patch seemingly simple boundary-case bugs, leaving many programs one step short of success.
-Experiments with stronger LLMs such as Qwen2.5-32B-Instruct and DeepSeek-V3 show that these larger LLMs seldom make such mistakes, indicating that their trajectories can serve as high-quality supervision for smaller LLMs.
-Challenges.
-Effective role-specific fine-tuning is hampered by two data gaps.
-(i) Public code-generation benchmarks provide only problem–solution pairs and contain no retrieval queries, step-wise plans, or debug traces; thus, no off-the-shelf corpus targets the intermediate process of the multi-agent coding pipeline.
-(ii) Even when synthetic trajectories are generated, many are noisy and fail the unit tests, so naïvely training on them would propagate errors rather than correct them.
-Methods.
 Fig.
 3 shows the proposed trajectory construction method.
-Our data pipeline begins by asking strong LLMs—Qwen2.5-32B-Instruct or DeepSeek-V3—to solve each coding task while explicitly printing the intermediate artefacts (i.e.
-trajectories) obtained from each MapCoder role.
-In one shot, the model emits an algorithm description for the retrieval agent, a step-wise plan for the planning agent, source code for the coding agent, and a patched version for the debugging agent.
-Because the same LLM generates the entire trajectory, the four outputs are mutually consistent and naturally aligned with their respective prompts.
-To convert these raw trajectories into reliable supervision, we exploit the fact that MapCoder’s final objective—passing all the unit tests—is objectively verifiable.
-After executing the LLM-produced code, we keep the complete trajectory only if the program passes.
-This pass‐based filtering eliminates noisy or partially correct traces without any manual inspection: a failure at test time almost always implies that at least one upstream artefact (faulty retrieval cue, flawed plan, or ineffective patch) is wrong, so discarding the whole sequence prevents error propagation during fine-tuning.
-The surviving set—about 2,200 retrieval examples and 4,200 debugging examples from APPS, xCodeEval, and CodeContests—forms a high-quality, role-aligned corpus that directly trains rank-32 LoRA on the shared 7B backbone.
-Empirically, models tuned on this filtered dataset regain XML-schema fidelity and bug-repair accuracy comparable to the original 32 B system while retaining the memory footprint advantages of a 7B model (see Table 2 for resulting accuracy enhancement).
+Our data pipeline begins by asking strong LLMs, namely Qwen2.5-32B-Instruct and DeepSeek-V3, to solve each coding task while explicitly printing intermediate artefacts (i.e., trajectories) produced by each MapCoder role.
+To build reliable training data, we collect trajectories from strong LLMs and keep only those whose final code passes all unit tests.
+Unlike rejection sampling based solely on local validity at the single-agent level Zelikman et al.
+(2022); Zeng et al.
+(2023), our pass-based filtering ensures end-to-end success across all roles.
+This focuses fine-tuning on trajectories verified through full execution, yielding accuracy gains over locally valid samples (Table 1, row 2 vs.
+row 3).
+Retrieval Planning / Coding xCodeEval
+Dataset Filtering Dataset Filtering Accuracy (%)
+11.32
+Strong 16.04
+Strong Format 16.98
+Strong Pass Strong Pass 18.87
+Strong Pass Supervisor Pass 22.64
+Table 1: Ablation study on data source and filtering methods for training retrieval, planning, and coding agents on xCodeEval.
+“Format” filtering keeps samples with valid single-agent outputs, while “Pass” retains only those where the final program passes all unit tests.
+“Strong” denotes data from a large LLM; “Supervisor” indicates trajectories refined after failure analysis.
+Figure 4: Cosine similarity between algorithm descriptions generated by base and fine-tuned models and ground-truth algorithm tags for 50 xCodeEval problems. Darker cells indicate stronger alignment.
+Our analysis further shows that self-collected trajectories from a 7B model often mislabel tasks, such as overpredicting dynamic programming, and exhibit limited algorithmic diversity.
+As a result, fine-tuning on these traces leads to lower performance (7BFT).
+In contrast, trajectories generated by 32B align more closely with ground-truth tags and provide more accurate algorithm descriptions.
+Fine-tuning on these trajectories (7BFT(32B)) significantly improves performance, demonstrating that strong-LLM supervision transfers both correctness and algorithmic diversity to smaller models.
 Figure 5: Supervisor-aided data collection pipeline.
@@
 If the revised result passes, the updated trajectory is added to the fine-tuning dataset.
-Retrieval Planning / Coding xCodeEval
-Dataset
-Pass
-Filtering
-Dataset
-Pass
-Filtering
-Accuracy (%)
-11.32
-Strong 16.04
-Strong 16.98
-Strong Strong 18.87
-Strong Supervisor 22.64
-Table 2: Ablation study on data source and pass-based filtering for training retrieval, planning, and coding agents on xCodeEval.
-“Strong” refers to data generated by a large LM (e.g., Qwen2.5-32B, DeepSeek-V3), and “Supervisor” denotes data refined by a supervisor model after failure analysis.
 5.2 Supervisor-Guided Planning and Coding
-Observations.
-Even after the role-specific fine-tuning on strong LLM trajectories, the 7B planning agent still emits fragmentary execution flows, omitting edge cases, misordering steps, or drifting from the prescribed outline.
-Also, the coding agent continues to generate programs that compile yet collapse under the full unit-test suite.
-These local errors propagate: an incomplete plan starves the coder of crucial hints, and a superficially correct program deceives the debugger until final evaluation.
-Table 2 confirms that this trajectory-only adaptation closes part of the gap but suffers from a noticeable accuracy loss.
-In contrast, Qwen2.5-32B-Instruct and DeepSeek-V3 produce structurally sound plans and largely correct first-pass implementations, indicating that stronger models possess global cues that the smaller model has not yet internalised.
-Challenges.
-Despite the availability of agent-wise strong-model trajectories, two obstacles remain for high-quality fine-tuning.
-First, a locally tuned agent lacks global awareness: improving its own output does not guarantee end-to-end success unless that output is judged in context.
-Second, the capacity gap hinders direct imitation: a 7B model exposed to large-scale models’ plans and coding results may underfit or copy surface patterns without learning the deeper reasoning (phenomenon called capacity gap Bansal et al.
+Even when fine-tuned on strong-model trajectories, 7B agents struggle to achieve robust multi-stage reasoning.
+One challenge is the lack of global awareness, where agents trained in isolation fail to account for downstream dependencies, an issue previously discussed in Section 4.
+Another is the capacity gap Bansal et al.
 (2024); Xu et al.
-(2025)).
-Methods.
-We propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model.
+(2025): while strong LLMs produce coherent and mostly correct outputs, small models tend to mimic surface forms without learning the underlying reasoning.
+These limitations underscore the need for supervision that supports both global coordination and deeper abstraction.
+To address these issues, we propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model.
 As shown in Fig.
-4, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory.
+5, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory.
 If the resulting program fails its unit tests, the entire trajectory is forwarded to a high-capacity supervisor LLM (DeepSeek-V3).
-The supervisor analyses the trace, pinpoints the agent primarily responsible for the failure, and issues concise, role-specific feedback, and only that agent regenerates its output.
+The supervisor analyzes the trace, identifies the agent primarily responsible for the failure, and issues concise, role-specific feedback.
+Only the selected agent then regenerates its output.
 The revised trajectory is re-tested; once all tests pass, the final plan–code pair is added to the fine-tuning corpus.
@@
 The supervisor operates only during data generation, keeping inference lightweight.
-We store just the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime.
+We store only the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime.
 Crucially, every example in this corpus is produced and execution-validated by the 7B model, eliminating concerns about a capacity gap.
-Iterating over thousands of problems yields a dataset that aligns planners and coders with end-to-end success.
+Iterating over thousands of problems yields a dataset that better aligns planners and coders with end-to-end success.
 Table 1 confirms that this strategy delivers a substantial accuracy boost.
-FT Method
-Trainable
-Parameters
-Required Memory
-for Training (GB)
-Accuracy
+5.3 Multi-Agent LoRA Fine-Tuning
+FT Method Trainable Parameters Required Memory for Training (GB) Accuracy (%)
 FFT 7615.62M 45.69 18.87
@@
 Accuracy is measured only up to the coding stage, without debugging.
-3.3 Multi-Agent LoRA Fine-Tuning
-To enable role-specific specialization in our multi-agent framework, we adopt an agent-wise LoRA approach.
-All four agents—retrieval, planning, coding, and debugging—share a frozen 7B base model, each paired with its own lightweight LoRA.
-These adapters are fine-tuned independently using pass-based-filtered and supervisor-guided trajectory distillation.
-This setup significantly reduces GPU memory usage during fine-tuning and supports modular training and deployment, allowing each agent to be updated without duplicating the full model.
-As shown in Table 3, LoRA significantly reduces the number of trainable parameters and memory footprint, while achieving even higher accuracy than full fine-tuning.
-We attribute this accuracy gain in part to LoRA’s implicit regularization, which steers the model toward role-specific behaviors without overwriting core pretrained knowledge—a finding consistent with prior work Hu et al.
-(2021); Houlsby et al.
-(2019).
-4 Experimental Settings
+Figure 6: Training and validation loss for the Planning agent using full fine-tuning (FFT) and LoRA.
+LoRA shows higher training loss but maintains lower and more stable validation loss, suggesting better generalization.
+To enable role-specific specialization, we adopt an agent-wise LoRA strategy: all four agents share a frozen 7B backbone, each with independent rank-32 LoRA adapters fine-tuned on role-specific data.
+Unlike prior approaches that fine-tune the entire model across tasks Shen et al.
+(2024); Zeng et al.
+(2023), our method isolates agent behavior at the adapter level while preserving a unified backbone.
 Method xCodeEval (106) APPS (150) CodeContests (165)
@@
 The numbers in parentheses indicate the number of problems in each benchmark.
+We empirically show that LoRA serves effectively as a modular control layer in multi-agent code generation, delivering both accuracy and efficiency gains (Table 2).
+We attribute this to LoRA’s implicit regularization Hu et al.
+(2021); Houlsby et al.
+(2019), which promotes task-specific behavior without overwriting pretrained knowledge.
+In multi-agent settings with complex, structured I/O, full fine-tuning often overfits to superficial patterns, weakening general reasoning.
+As shown in Fig.
+6, full fine-tuning achieves lower training loss but increasing validation loss, whereas LoRA maintains lower and more stable validation loss, indicating better generalization and adaptability across roles.
+6 Experimental Settings
 Models.
 All agent-wise fine-tuning experiments use Qwen2.5-7B-Instruct as the base model.
-For comparison, we include larger models—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, serving as upper bounds to assess the effectiveness of our lightweight strategy.
+For comparison, we include larger variants—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct—which serve as upper bounds for evaluating the effectiveness of our lightweight strategy.
+To assess whether a code-pretrained backbone improves performance, we additionally test Qwen2.5-Coder-7B-Instruct.
+We also apply our method to Llama3.1-8B and CodeLlama-7B to demonstrate generalizability beyond the Qwen family.
 Benchmarks.
@@
 For CodeContests, we evaluate all 165 official test problems.
-We randomly select 150 test-set problems for APPS, following MapCoder Islam et al.
+For APPS, we randomly select 150 test-set problems, following MapCoder Islam et al.
 (2024).
-Each benchmark includes natural language problem descriptions with sample I/O and is evaluated using hidden unit tests.
-We also optionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
-Training Configuration.
-Each agent is fine-tuned using LoRA adapters with rank 32 applied to the linear modules for query, key, value, and output projection.
-Training is implemented with Hugging Face Transformers and PEFT libraries HuggingFace , using a learning rate of 2e-5, gradient accumulation of 16, and 3 training epochs.
-Hyperparameters were tuned individually per agent to ensure stable and consistent performance.
-All training experiments were conducted on a single NVIDIA A100 80GB GPU.
-Evaluation Setting.
-We report pass@1 accuracy as the primary metric, reflecting whether the model’s top-1 output passes all test cases.
-All outputs are generated using greedy decoding for reproducibility and consistency with baseline evaluation.
-5 Experimental Results
+Each benchmark provides natural language problem descriptions with sample I/O and is evaluated using hidden unit tests.
+We additionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
+Fine-Tuning Setup.
+Each agent is fine-tuned using LoRA adapters with rank-32, applied to the linear layers for query, key, value, and output projections.
+We use a learning rate of 2e-5, gradient accumulation of 16, and train for three epochs, with hyperparameters tuned per agent to ensure stability.
+All training runs are conducted on a single NVIDIA A100 80GB GPU.
+The cost of data curation prior to fine-tuning is summarized in Table 8 (Section 7.5).
+Evaluation Setting. We report pass@1 accuracy using greedy decoding to ensure reproducibility and align with baseline evaluations.
+7 Results
 7.1 Effectiveness of Agent-wise Fine-Tuning
-Table 4 compares the performance of a small LM (Qwen2.5-7B) across various single-agent prompt optimization methods—direct prompting, chain-of-thought (CoT) Wei et al.
+Table 3 compares the performance of Qwen2.5-7B under various prompting strategies—direct prompting, CoTWei et al.
 (2023), self-planningJiang et al.
-(2024b), and analogical prompting Yasunaga et al.
-(2024)—as well as multi-agent approaches: MapCoder and our proposed MapCoder-Lite.
-Without fine-tuning, MapCoder shows only marginal gains over single-agent methods and even underperforms Direct prompting on xCodeEval (13.21% vs.
-18.87%), despite using four specialized agents.
-This highlights the limitations of small LMs in multi-agent setups, where poor coordination and frequent format violations (as indicated by numerous format failures) prevent reliable end-to-end performance.
-In contrast, MapCoder-Lite, with agent-wise fine-tuning, achieves substantial improvements across all benchmarks.
+(2024b), analogical Yasunaga et al.
+(2024)—and multi-agent setups.
+Without fine-tuning, MapCoder yields only marginal improvements and even underperforms direct prompting on xCodeEval (13.21% vs.
+18.87%), underscoring the difficulty of coordinating small models in multi-agent pipelines due to format errors and poor inter-agent coordination.
+In contrast, MapCoder-Lite with agent-wise fine-tuning achieves substantial improvements across all benchmarks.
 On xCodeEval, accuracy more than doubles to 28.30%, outperforming all prompting baselines.
-Similar gains are seen on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%).
+Similar gains are observed on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%).
 Notably, fine-tuning eliminates all format failures, indicating that even small LMs can produce structurally consistent outputs when role-specialized.
-These results emphasize the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks.
+These results highlight the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks, with qualitative examples in Appendix E further illustrating improvements across all roles.
 7.2 Comparison with Larger Models
@@
 Table 4: Pass@1 accuracy (%) of MapCoder using Qwen models of different scales on xCodeEval, APPS, and CodeContests. Memory usage (GB) and Time Per Output Tokens (ms).
+7.3 Ablation Study
+Contribution of Individual Agent Fine-Tuning.
+Table 5 presents an ablation study of agent-wise fine-tuning on xCodeEval.
+Tuning the retrieval agent alone raises accuracy from 13.21% to 18.87% and cuts format failures from 29 to 12.
+Adding the planning agent further improves accuracy to 26.42% and eliminates all format errors, boosting both non-debug and debug-assisted completions.
+Fine-tuning the coding agent (without debugging) increases non-debug passes to 24 but reduces debug-assisted ones, as stronger initial code often bypasses the debugger.
+In contrast, tuning the debugging agent (without coding) improves recovery from weak code, achieving 27.36% accuracy and 12 debug-assisted completions.
+Full fine-tuning of all agents yields the best performance (28.30%), confirming that each agent contributes uniquely and coordinated tuning is critical for optimal results.
 Agent xCodeEval Pass Pass Format
@@
 Table 5: Ablation study of per-agent fine-tuning contributions in MapCoder on xCodeEval. R, P, C, and D denote Retrieval, Planning, Coding, and Debugging agents, respectively.
+Retrieval Planning Coding xCodeEval (%)
+32 32 32 22.64
+8 32 32 18.87
+32 8 32 20.75
+32 32 8 21.70
+8 8 8 21.70
+64 32 32 16.98
+32 64 32 20.75
+32 32 64 20.75
+64 64 64 16.98
+Table 6: Effect of LoRA rank on multi-agent performance. The three leftmost columns indicate the LoRA ranks applied to the retrieval, planning, and coding agents, respectively.
 Method HumanEval (164 problems) MBPP (397 problems)
@@
 Table 7: Evaluation of MapCoder and MapCoder-Lite on HumanEval and MBPP benchmarks.
-5.3 Ablation Study
-Contribution of Individual Agent Fine-Tuning.
-Table 6 shows an ablation study evaluating the impact of fine-tuning each agent in MapCoder on xCodeEval.
-Fine-tuning the retrieval agent alone improves accuracy from 13.21% to 18.87% and reduces format failures from 29 to 12, indicating better structural consistency.
-Adding the planning agent further boosts accuracy to 26.42% and eliminates format errors entirely, increasing both non-debug passes (17 → 18) and debug-assisted completions (3 → 10).
-Fine-tuning the coding agent (without debugging) raises non-debug passes to 24 but reduces debug-assisted passes to 2, since MapCoder only invokes the debugger when initial code fails sample I/O tests.
-Stronger code may bypass debugging entirely, leaving some errors uncaught.
-Conversely, fine-tuning the debugging agent (without coding) increases accuracy to 27.36%, with 12 debug-assisted completions—showing that an effective debugger can recover from weaker initial code.
-When all four agents are fine-tuned, overall accuracy peaks at 28.30%, with a strong balance between direct (24) and debug-assisted (6) completions.
-These results demonstrate that each agent addresses distinct failure modes and that coordinated fine-tuning is essential for maximizing end-to-end performance.
+Impact of LoRA Rank.
+We study the effect of LoRA rank on multi-agent performance to justify the configuration used in this work.
+The rank-32 setting was selected based on empirical evaluation rather than by convention.
+As shown in Table 6, rank-32 achieves the highest xCodeEval accuracy (22.64%) among all tested configurations.
+Lower-rank settings (e.g., $(8,8,8)$) underperform despite being parameter-efficient, while higher-rank configurations (e.g., 64) do not yield further gains and can degrade performance.
+These results suggest that multi-agent performance depends on interactions between agent roles rather than parameter count alone.
+Overall, rank-32 offers the best balance between accuracy and efficiency, introducing only about 3% additional trainable parameters relative to the backbone.
+Backbone Selection.
+Choosing the right backbone is critical for multi-agent code generation, where each agent performs a distinct, reasoning-intensive role.
+Among SLM candidates, we compare a general-purpose model (Qwen2.5-7B-Instruct Qwen et al.
+(2025)) and a code-specialized variant (Qwen2.5-Coder-7B-Instruct Hui et al.
+(2024)) for use in MapCoder.
+Despite similar sizes, the general-purpose model achieves higher accuracy and format adherence.
+Agent-wise ablations further show that replacing even one agent with its coder counterpart degrades performance, suggesting that the coder model lacks the contextual reasoning and adaptability needed for role-specific tasks.
+These results support our choice of the general-purpose model as the unified backbone for MapCoder-Lite (see Appendix B).
+7.4 Generalization and Robustness
 Evaluation on Simpler Coding Tasks.
@@
 These results suggest that MapCoder-Lite not only preserves generalization capability but also improves robustness on structurally simpler tasks.
-Multi-Agent Tuning with Coder Model.
-We evaluate whether multi-agent code generation can be enhanced using a code-specialized language model (a.k.a.
-Coder model), such as Qwen2.5-Coder-7B-Instruct Hui et al.
-(2024).
-As shown in Table 8, applying the multi-agent framework (MapCoder) and our fine-tuning strategy (MapCoder-Lite) to the Coder model does improve its performance, demonstrating that even domain-specialized models benefit from structured, role-specific tuning.
-However, despite these gains, the Coder-based pipeline still lags significantly behind our baseline model, Qwen2.5-7B-Instruct, both before and after fine-tuning.
-We attribute this gap to the Coder model’s limited general reasoning ability, which is critical for handling the diverse roles in multi-agent workflows, including retrieval, planning, and debugging, in addition to coding.
-In contrast, the general-purpose Qwen2.5-7B-Instruct model more effectively supports multi-agent coordination, and with our memory-efficient fine-tuning (MapCoder-Lite), achieves even higher end-to-end accuracy.
-Model Method Accuracy (%)
-Qwen2.5-Coder-7B-Instruct MapCoder 5.66
-MapCoder-Lite 15.09
-Qwen2.5-7B-Instruct MapCoder 11.32
-MapCoder-Lite 22.64
-Table 8: Comparison of MapCoder and MapCoder Lite using Qwen2.5-Coder-7B-Instruct. Accuracy is measured only up to the coding stage, without debugging.
-5.4 Case Study
-To qualitatively assess the impact of agent-wise fine-tuning, we examine representative before-and-after examples for each agent.
-As illustrated in Appendix B, fine-tuning leads to notable improvements in algorithm retrieval, plan coverage, code correctness, and bug resolution.
-These examples highlight how each agent benefits from role-specific supervision, supporting the effectiveness of our fine-tuning strategy.
+Model Compatibility.
+Our method is architecture-agnostic and applicable to LoRA-supported backbones.
+Beyond the Qwen2.5 family, MapCoder-Lite improves accuracy and reduces format failures on Llama3.1-8B and CodeLlama-7B across xCodeEval and HumanEval (Appendix D).
+We also evaluate MapCoder-Lite on the recent reasoning-oriented model Qwen3-4B; despite strong single-agent prompting performance, stable multi-agent behavior is only achieved after agent-wise fine-tuning, with detailed results reported in Appendix C.
+7.5 Data Curation Cost and Statistics
+Table 8 summarizes the dataset size, resource usage, and data generation time for each agent.
+In particular, we curate approximately 2.3k execution-verified trajectories for the retrieval agent, 1.1k each for the planning and coding agents, and 4.2k for the debugging agent.
+We used 2×A100 GPUs with vLLM serving to run Qwen2.5-32B and Qwen2.5-7B models, and partially relied on the DeepSeek API as supervisor and debugging agent, incurring \$20–30.
+Data generation is performed only once, and the resulting datasets are reusable across models and experiments.
+The strong supervisor model is invoked only in failure cases to provide brief corrective feedback, rather than full trajectory generation, keeping strong-model involvement infrequent and tightly bounded to training-time supervision.
+Retrieval and debugging took longer due to longer outputs; debugging in particular depends on planning and coding outputs, requiring upstream completion.
+While data was generated per agent during development, the pipeline can be streamlined into a single pass for faster future use.
+Retrieval Planning, Coding Debugging
+Dataset size 2.3k 1.1k (each) 4.2k
+Time 2–3 days 16 hours 4–5 days
+GPUs 2 $\times$ A100 2 $\times$ A100 2 $\times$ A100
+Table 8: Dataset size, data generation time, and computational resource usage for data curation for each agent.
 8 Conclusion
 We introduced MapCoder-Lite, a multi-agent coding framework built on a 7B model with lightweight LoRA adapters.
-Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance with lower cost.
-Our results show that small models, when carefully fine-tuned, can match the reliability and accuracy of larger systems.
+Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance at substantially lower cost.
+Our results demonstrate that small language models, when carefully fine-tuned, can match the reliability and accuracy of much larger systems in multi-agent code generation.
+Acknowledgments
+This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2025-00561961).
+This work was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.
+RS-2025-02214497, Development of Low-Level Optimization Program API Technology for AI Semiconductors).
 Limitations
-Our framework relies on distilled trajectories from strong models like DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance.
-Rather than indicating a ceiling, this gap highlights opportunities to further enhance small models—potentially through architectural extensions, better training objectives, or reinforcement-based learning from distillation feedback.
-In addition, the multi-agent structure introduces a wide design space for fine-tuning, as each agent has a distinct role and optimization goal.
-While we adopted a uniform tuning recipe across agents, this modularity also enables targeted specialization and co-training strategies in future work, suggesting that more adaptive tuning could unlock even greater performance.
+Our framework relies on distilled trajectories from strong models such as DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance.
+Rather than indicating a hard performance ceiling, this gap highlights opportunities to further enhance small models, potentially through architectural extensions or improved training objectives.
+In addition, the multi-agent structure introduces a broad design space for fine-tuning, as each agent has a distinct role and optimization objective.
+While we adopt a uniform tuning recipe across agents in this work, more adaptive tuning could further improve performance.
 Appendix A Details of Supervisor
@@
 7 and Fig.
-6 illustrate the operation of the supervisor used in our data curation pipeline.
-The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, plan, code), and the result of test case execution.
+8 illustrate the operation of the supervisor in our data curation pipeline.
+The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, step-by-step plan, and code), and the result of test case execution.
 As shown in Fig.
-5, the input prompt explicitly instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent.
-The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that can be used to re-generate only the faulty component.
+7, the input prompt instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent.
+The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that enables re-generation only the faulty component.
 Fig.
-6 shows a typical response from the supervisor.
+8 shows a representative response from the supervisor.
 It correctly attributes the failure to the planning agent, explains why the plan is insufficient, and suggests how it should be modified.
-This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without requiring human annotation.
-Appendix B Improvements After Fine-Tuning
-Fig. 789, and 10 illustrate representative improvements observed in the 7B model for the retrieval, planning, coding, and debugging agents, respectively.
+This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without human annotation.
+Appendix B Backbone Model Comparison: General-Purpose vs. Coder
+To investigate the impact of backbone model selection in multi-agent code generation, we conducted a series of controlled experiments comparing general-purpose and code-specialized models.
+Specifically, we tested Qwen2.5-7B-Instruct (general-purpose) and Qwen2.5-Coder-7B-Instruct (code-specialized) under various configurations of the MapCoder and MapCoder-Lite pipelines.
+The results show that general-purpose models are more robust across all agent roles and better suited for both zero-shot and fine-tuned multi-agent setups.
+Benchmark Performance (Zero-shot).
+Table 9 shows the performance of MapCoder using each backbone model.
+The general-purpose variant consistently achieves higher accuracy and fewer format failures, highlighting its superiority in multi-agent coordination and instruction-following tasks.
+Benchmark Metric 7B 7BCoder
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 9: Comparison of Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct across benchmarks.
+Agent-wise Ablation (Zero-shot).
+As shown in Table 10, replacing any agent in the general-purpose pipeline with its coder counterpart leads to a consistent accuracy drop, most notably in the planning and coding agents.
+Even for the coding role, where the coder model is expected to excel, performance decreases—highlighting the importance of upstream integration and general reasoning.
+Retrieval Planning Coding Debugging xCodeEval (%)
+7B 7B 7B 7B 13.21
+7BCoder 7B 7B 7B 12.26
+7B 7BCoder 7B 7B 10.38
+7B 7B 7BCoder 7B 10.38
+7B 7B 7B 7BCoder 12.26
+Table 10: Ablation of 7B vs. 7B Coder in retrieval, planning, coding, and debugging agents on the xCodeEval benchmark.
+Training Loss after Fine-tuning.
+We fine-tuned both models using identical LoRA settings across all agents.
+As summarized in Table 11, the coder model converges to higher final loss values, suggesting a poorer fit to role-specialized training data.
+This pattern is consistent across all agents.
+Model Retrieval Planning Coding Debugging
+Qwen2.5-7B-Instruct 0.34 0.25 0.08 0.33
+Qwen2.5-Coder-7B-Instruct 0.36 0.35 0.13 0.34
+Table 11: Final training loss after LoRA fine-tuning for each agent using general-purpose vs. coder-specific models.
+Ablation after LoRA Fine-tuning.
+Table 12 presents accuracy when substituting each LoRA-finetuned agent with its coder-based variant.
+Accuracy declines in all settings, and the fully coder-based pipeline achieves only 18.87%, compared to 28.30% for the general-purpose variant.
+These results indicate that even after fine-tuning, general-purpose models better support the multi-agent workflow.
+Retrieval Planning Coding Debugging xCodeEval (%)
+7BFT 7BFT 7BFT 7BFT 28.30
+7BCoderFT 7BFT 7BFT 7BFT 21.70
+7BFT 7BCoderFT 7BFT 7BFT 23.58
+7BFT 7BFT 7BCoderFT 7BFT 24.53
+7BFT 7BFT 7BFT 7BCoderFT 23.58
+7BCoderFT 7BCoderFT 7BCoderFT 7BCoderFT 18.87
+Table 12: xCodeEval accuracy when substituting each agent with a LoRA-finetuned coder-specific model (7BCoderFT), compared against the general-purpose baseline (7BFT).
+Appendix C Evaluation on Recent Reasoning Models (Qwen3)
+We evaluate MapCoder-Lite on Qwen3-4B, a recently released small language model that significantly improves single-agent reasoning and outperforms Qwen2.5-7B under direct prompting.
+As shown in Table 13, Qwen3-4B achieves strong direct-prompting performance (22.64% without thinking and 26.42% with thinking).
+However, this improvement does not directly translate into reliable multi-agent behavior.
+When deployed in the MapCoder pipeline without fine-tuning, Qwen3-4B attains only 3.77% accuracy, exhibiting the same instability observed with earlier backbones.
+Applying MapCoder-Lite mitigates this issue.
+After fine-tuning the retrieval, planning, and coding agents, Qwen3-4B achieves 30.19% accuracy, surpassing the fine-tuned Qwen2.5-7B despite using fewer parameters.
+These results indicate that recent advances in single-agent reasoning alone are insufficient for stable multi-agent workflows, and that targeted agent-wise specialization remains essential.
+Moreover, the larger gains observed with Qwen3-4B suggest that MapCoder-Lite scales positively with backbone capability, effectively activating role-specific behaviors that are not induced by prompting alone.
+Method Qwen2.5-7B-Instruct Qwen3-4B
+Direct (non-thinking) 18.87 22.64
+Direct (thinking) 26.42
+MapCoder (RPC) 11.32 3.77
+MapCoder-Lite (RPC) 22.64 30.19
+Table 13: xCodeEval performance of Qwen2.5-7B and Qwen3-4B under direct prompting and MapCoder pipelines. RPC denotes the retrieval–planning–coding agents.
+Appendix D Generalization Across Model Families
+While our main experiments used the Qwen series due to their strong coding performance on HumanEval and MBPP during development, our method remains architecture-agnostic.
+Since MapCoder-Lite applies LoRA through the PEFT library, it can be seamlessly used with any compatible model.
+To test this generalizability, we evaluated both MapCoder and MapCoder-Lite using Llama3.1-8B-Instruct and CodeLlama-7B-Instruct-hf as backbones.
+As shown in Table 14, both models performed poorly without fine-tuning on competitive programming tasks.
+For example, Llama3.1-8B scored 0.00% with 94 format failures on xCodeEval.
+After applying MapCoder-Lite, Llama3.1-8B achieved 16.04% accuracy with zero format failures, demonstrating the adaptability of our pipeline.
+In contrast, CodeLlama-7B showed limited improvement on xCodeEval but responded well on function-level benchmarks such as HumanEval.
+Model Benchmark MapCoder MapCoder-Lite
+Accuracy (%) Accuracy (%)
+Llama3.1-8B-Instruct xCodeEval 0.00 16.04
+CodeLlama-7B-Instruct xCodeEval 0.94 2.83
+CodeLlama-7B-Instruct HumanEval 10.98 45.12
+Table 14: Comparison of MapCoder and MapCoder-Lite on Llama3.1-8B and CodeLlama-7B models on xCodeEval and HumanEval benchmarks.
+Additionally, we compare MapCoder-Lite with various prompting strategies using the Llama3.1-8B backbone.
+As summarized in Table 15, MapCoder underperforms direct prompting baselines, whereas MapCoder-Lite substantially improves accuracy and surpasses all prompting methods.
+These results reinforce that multi-agent pipelines require targeted fine-tuning to be effective: without it, they underperform simpler approaches, while with it, they become competitive and scalable.
+Method Accuracy (%)
+Direct 6.60
+CoT 4.72
+Self-planning 4.72
+Analogical 5.66
+Multi-agent (MapCoder) 0.00
+Multi-agent + FT (MapCoder-Lite) 16.04
+Table 15: Accuracy comparison between MapCoder-Lite and other prompting baselines on xCodeEval using Llama3.1-8B-Instruct.
+Appendix E Improvements After Fine-Tuning
+Fig. 91011, and 12 illustrate representative improvements observed in the 7B model for the retrieval, planning, coding, and debugging agents, respectively.
 In the retrieval example (Fig.
-7), the pre-fine-tuned model incorrectly identifies the core algorithm as “Sliding Window” and includes unsupported tags like <description> within the <algorithm> block, failing to conform to the expected XML schema.
+9), the pre-fine-tuned model incorrectly identifies the core algorithm as “Sliding Window” and includes unsupported tags like <description> within the <algorithm> block, failing to conform to the expected XML schema.
 After fine-tuning, the agent accurately identifies the algorithm (“Counting and Matching Pairs”) and outputs a well-structured XML response with all required tags properly closed.
 In the planning example (Fig.
-8), the original plan fails to capture key logical conditions (e.g., that the input must be both even and greater than 2).
+10), the original plan fails to capture key logical conditions (e.g., that the input must be both even and greater than 2).
 After fine-tuning, the agent successfully generates a plan that handles these conditions explicitly and correctly guides the coding agent.
 In the coding example (Fig.
-9), the unfine-tuned model incorrectly processes multiple input values on a single line, violating the problem’s input specification.
+11), the unfine-tuned model incorrectly processes multiple input values on a single line, violating the problem’s input specification.
 Fine-tuning enables the agent to read inputs line by line, aligning its behavior with the expected input format and improving functional correctness.
 Finally, the debugging example (Fig.
-10) demonstrates how the unfine-tuned model fails to fix a bug caused by improper input parsing.
+12) demonstrates how the unfine-tuned model fails to fix a bug caused by improper input parsing.
 The fine-tuned debugging agent correctly diagnoses the issue and proposes a revised plan and code that successfully passes all test cases.
-These examples highlight the role-specific benefits of fine-tuning, improving not only task-level correctness but also structural and logical alignment across the entire pipeline.
+Appendix F Trajectory Example
+These are example trajectories used for fine-tuning each agent. Each consists of an input–output pair, and the examples are drawn from different problems.
+F.1 Retrieval Agent
+Example Trajectory of Retrieval (Input) Given a problem, algorithm tags, and solution code, provide relevant problems then identify the algorithm behind it and also explain the tutorial of the algorithm.
+# Problem: You are given several queries.
+Each query consists of three integers $p$, $q$ and $b$.
+You need to answer whether the result of $p/q$ in notation with base $b$ is a finite fraction.
+A fraction in notation with base $b$ is finite if it contains finite number of numerals after the decimal point.
+It is also possible that a fraction has zero numerals after the decimal point.
+—–Input—– The first line contains a single integer $n$ ($1\leq n\leq 10^{5}$) — the number of queries.
+Next $n$ lines contain queries, one per line.
+Each line contains three integers $p$, $q$, and $b$ ($0\leq p\leq 10^{18}$, $1\leq q\leq 10^{18}$, $2\leq b\leq 10^{18}$).
+All numbers are given in notation with base 10.
+—–Output—– For each question, in a separate line, print Finite if the fraction is finite and Infinite otherwise.
+—–Examples—– Input 2 6 12 10 4 3 10 Output Finite Infinite Input 4 1 1 2 9 36 2 4 12 3 3 5 4 Output Finite Finite Finite Infinite —–Note—– $\frac{6}{12}=\frac{1}{2}=0.5_{10}$ $\frac{4}{3}=1.(3)_{10}$ $\frac{9}{36}=\frac{1}{4}=0.01_{2}$ $\frac{4}{12}=\frac{1}{3}=0.1_{3}$ Sample Input Format: 2 6 12 10 4 3 10 Sample Output Format: Finite Infinite ——- Important: You must follow the input output format.
+Input should be taken from standard input and output should be given to standard output.
+Note: If you are writing a function then after the function definition take input from using input() function, call the function with specified parameters and finally print the output of the function.
+# Exemplars: Recall three (03) relevant and distinct problems (different from problem mentioned above).
+For each problem, 1.
+describe it 2.
+generate Python3 code step by step to solve that problem 3.
+finally generate a planning to solve that problem # Algorithm: —————- Important: Your response must follow the following xml format- <root> <problem> # Recall three (03) relevant and distinct problems (different from problem mentioned above).
+Write each problem in the following format.
+<description> # Describe the problem.
+</description> <code> # Let’s think step by step to solve this problem in Python3 programming language.
+</code> <planning> # Planning to solve this problem.
+</planning> </problem> # similarly add more problems here… <algorithm> # Identify the algorithm (Brute-force, Dynamic Programming, Divide-and-conquer, Greedy, Backtracking, Recursive, Binary search, and so on) that needs to be used to solve the original problem.
+# Write a useful tutorial about the above mentioned algorithms.
+Provide a high level generic tutorial for solving this types of problem.
+Do not generate code.
+</algorithm> </root>
+Example Trajectory of Retrieval (Output) <root> <problem> <description> Given a fraction $p/q$, determine if it can be represented as a finite decimal in base $b$.
+The fraction is finite in base $b$ if the denominator $q$ (after simplifying the fraction) has no prime factors other than those of the base $b$.
+</description> <code> import math def solve(): n = int(input()) for _ in range(n): p, q, b = map(int, input().split()) # Simplify the fraction p/q g = math.gcd(p, q) q_simplified = q // g # Remove all common factors between q_simplified and b while True: g = math.gcd(q_simplified, b) if g == 1: break q_simplified //= g if q_simplified == 1: print("Finite") else: print("Infinite") solve() </code> <planning> 1.
+Read the number of queries.
+2.
+For each query, read $p$, $q$, and $b$.
+3.
+Simplify the fraction $p/q$ by dividing both numerator and denominator by their greatest common divisor (GCD).
+4.
+Repeatedly divide the simplified denominator by the GCD of the denominator and the base $b$ until no more common factors exist.
+5.
+If the resulting denominator is 1, the fraction is finite in base $b$; otherwise, it is infinite.
+6.
+Print "Finite" or "Infinite" based on the result.
+</planning> </problem> (Other example problems…) <algorithm> The algorithm used to solve the original problem is based on number theory, specifically the properties of fractions and their representations in different bases.
+The key insight is that a fraction $p/q$ has a finite representation in base $b$ if and only if the denominator $q$ (after simplifying the fraction) has no prime factors other than those of the base $b$.
+Tutorial: 1.
+Simplify the fraction $p/q$ by dividing both numerator and denominator by their GCD.
+2.
+Remove all common prime factors between the simplified denominator and the base $b$ by repeatedly dividing the denominator by the GCD of the denominator and the base.
+3.
+If the resulting denominator is 1, the fraction is finite in base $b$; otherwise, it is infinite.
+This approach efficiently checks the finiteness of the fraction by leveraging GCD operations to eliminate common factors, ensuring optimal performance even for large inputs.
+</algorithm> </root>
+F.2 Planning Agent
+Example Trajectory of Planning (Input) Given a competitive programming problem generate a concrete planning to solve the problem.
+# Problem: Given a string, reverse each word in the string while keeping the order of the words the same.
+For example, "Hello World" becomes "olleH dlroW".
+# Planning: 1.
+Split the input string into individual words using the ‘split()‘ method.
+2.
+Iterate over each word and reverse it using slicing (‘[::-1]‘).
+3.
+Join the reversed words back into a single string with spaces separating them.
+4.
+Print the result.
+## Relevant Algorithm to solve the next problem: The algorithm used to solve the original problem (extracting the first character of each word in a string) is a **String Manipulation** technique, specifically involving splitting the string into words and processing each word individually.
+### Tutorial: String Manipulation for Word Processing 1.
+**Splitting the String**: Use the ‘split()‘ method to divide the input string into a list of words based on spaces.
+2.
+**Processing Each Word**: Iterate over the list of words and extract the first character of each word.
+3.
+**Combining Results**: Join the extracted characters into a single string to form the final output.
+This approach is efficient for problems where you need to process individual words in a string.
+It leverages basic string operations and does not require complex algorithms like dynamic programming or backtracking.
+The key steps are splitting, iterating, and combining, which are fundamental in string manipulation tasks.
+## Problem to be solved: In this exercise, a string is passed to a method and a new string has to be returned with the first character of each word in the string.
+For example: “‘ "This Is A Test" ==> "TIAT" “‘ Sample Input Format: sees eyes xray yoat Sample Output Format: sexy ——- Important: You must follow the input output format.
+Input should be taken from standard input and output should be given to standard output.
+Note: If you are writing a function then after the function definition take input from using ‘input()‘ function, call the function with specified parameters and finally print the output of the function.
+## Sample Test cases: Input: sees eyes xray yoat Expected output: sexy Input: brown eyes are nice Expected output: bean ## Planning: —————- Important: You should give only the planning to solve the problem.
+Do not add extra explanation or words.
+Example Trajectory of Planning (Output) 1.
+Split the input string into individual words using the ‘split()‘ method.
+2.
+Iterate over each word and extract the first character.
+3.
+Join the extracted characters into a single string.
+4.
+Print the result.
+F.3 Coding Agent
+Example Trajectory of Coding (Input) Given a competitive programming problem generate Python3 code to solve the problem.
+## Relevant Algorithm to solve the next problem: The algorithm used in the original problem is **String Slicing**.
+### Tutorial: String Slicing String slicing is a technique to extract parts of a string by specifying start, end, and step indices.
+The general syntax is string[start:end:step], where: - start is the index where the slice begins (inclusive).
+- end is the index where the slice ends (exclusive).
+- step determines the stride (e.g., 1 for forward, -1 for reverse).
+#### Key Points: 1.
+**Indices**: Python uses zero-based indexing.
+Negative indices count from the end (1-1 is the last character).
+2.
+**Defaults**: If omitted, start defaults to 0, end to len(string), and step to 1.
+3.
+**Applications**: - Extracting substrings (s[1:4]).
+- Reversing strings (s[::-1]).
+- Removing characters (s[1:-1] removes first and last characters).
+#### High-Level Approach: 1.
+Identify the substring you need by determining the start and end indices.
+2.
+Use slicing to extract the substring.
+3.
+Handle edge cases (e.g., empty strings, out-of-bound indices).
+This technique is efficient ($O(k)$ for a slice of length $k$) and widely used in string manipulation tasks.
+## Problem to be solved: It’s pretty straightforward.
+Your goal is to create a function that removes the first and last characters of a string.
+You’re given one parameter, the original string.
+You don’t have to worry with strings with less than two characters.
+Sample Input Format: eloquent Sample Output Format: loquen ——- Important: You must follow the input output format.
+Input should be taken from standard input and output should be given to standard output.
+Note: If you are writing a function then after the function definition take input from using input() function, call the function with specified parameters and finally print the output of the function.
+## Planning: 1.
+Define a function remove_first_last_char that takes a string s as input.
+2.
+Use string slicing to exclude the first and last characters (s[1:-1]).
+3.
+Take input from the user, call the function, and print the result.
+## Sample Test cases: Input: eloquent Expected output: loquen Input: country Expected output: ountr ## Let’s think step by step.
+—————- Important: ## Your response must contain only the Python3 code to solve this problem.
+Do not add extra explanation or words.
+Example Trajectory of Coding (Output) “‘python
+def remove_first_last_char(s):
+  return s[1:-1]
+input_str = input()
+print(remove_first_last_char(input_str))
+F.4 Debugging Agent
+Example Trajectory of Debugging (Input) Given a competitive programming problem you have generated Python3 code to solve the problem.
+But the generated code can not pass sample test cases.
+Improve your code to solve the problem correctly.
+## Relevant Algorithm to solve the next problem: The algorithm used to solve the original problem of breaking up camel casing involves a straightforward approach to identifying capital letters within a string and inserting a space before them.
+This can be categorized under the Brute-force method since it directly iterates through the string and applies a simple condition to achieve the desired outcome.
+**Tutorial on Brute-force Algorithms:** Brute-force algorithms involve trying every possible solution until the correct one is found.
+They are often the simplest and most direct way to solve a problem but may not always be the most efficient, especially for larger datasets.
+The key aspects of implementing a brute-force algorithm include: - Identifying all possible solutions.
+- Checking each solution against the problem constraints.
+- Selecting the first solution that meets all criteria.
+In the context of the camel case problem, the brute-force approach directly iterates through the string, checks if a character is uppercase, and if so, inserts a space before it.
+This method is effective for the task at hand due to its simplicity and directness.
+## Problem to be solved: Complete the solution so that the function will break up camel casing, using a space between words.
+### Example “‘ solution("camelCasing") == "camel Casing" “‘ Sample Input Format: helloWorld Sample Output Format: hello World ——- Important: You must follow the input output format.
+Input should be taken from standard input and output should be given to standard output.
+Note: If you are writing a function then after the function definition take input from using input() function, call the function with specified parameters and finally print the output of the function.
+## Planning: 1.
+Identify the indices where a transition from lowercase to uppercase letter occurs.
+2.
+Insert a space before each identified index.
+3.
+Join the modified string segments.
+4.
+Return or print the final result.
+## Code: “‘ def solution(s): return ”.join([’ ’ + c if c.isupper() else c for c in s])[1:] s = input() print(solution(s)) “‘ ## Test Report: ## Tested passed: No test cases passed.
+## Tests failed: Input: helloWorld Expected Output: hello World Your Output: ello World Input: camelCase Expected Output: camel Case Your Output: amel Case ## Modified Planning: ## Let’s think step by step to modify Python3 Code for solving this problem.
+—————- Important: ## Note: Strictly follow the input and output format.
+The input should be taken from Standard input and output should be given to standard output.
+If you are writing a function then after the function definition take input using input() function then call the function with specified parameters and finally print the output of the function.
+Do not add extra print statement otherwise it will failed the test cases.
+## Your response must contain the modified planning and then the Python3 code inside “‘ block to solve this problem.
+Example Trajectory of Debugging (Output) ## Modified Planning: 1.
+**Identify Uppercase Letters**: Iterate through each character in the input string to check if it is uppercase.
+2.
+**Insert Space Before Uppercase Letters**: For each uppercase character found, insert a space before it.
+3.
+**Construct Result String**: Combine all characters, including the inserted spaces, to form the final result string.
+4.
+**Handle Edge Cases**: Ensure that the first character of the string is not modified if it is uppercase (no space should be inserted before the first character).
+The issue with the original code was that it was slicing the string from the second character ([1:]), which caused the first character to be omitted.
+This needs to be fixed by ensuring the first character is included in the result.
+## Modified Python3 Code: “‘python def solution(s): result = [] for c in s: if c.isupper(): result.append(’ ’ + c) else: result.append(c) return ”.join(result) s = input() print(solution(s)) “‘
 Figure 7: Prompt for the Supervisor.